# Integrable Kuralay equations: geometry, solutions and generalizations

### v3

@@ 6.2 Hirota bilinear form of the K-IE
$\epsilon^{1}$ ::: $ig_{1t}+g_{1xt}=0,$ (6.19)
$\epsilon^{3}$ ::: $\left[i\partial_{t}+\partial_{x}\partial_{t}\right]g_{3}=\left[iD_{t}-D_{x}D_{t}\right](\bar{f}_{2}.g_{1}),$ (6.20)
[…]
$\epsilon^{2n+1}$ ::: $\left[i\partial_{t}+\partial_{x}\partial_{t}\right]g_{2n+1}=\sum_{k+m=n}\left[iD_{t}-D_{x}D_{t}\right](\bar{f}_{2k}.g_{2m+1}),$ (6.24)
@@
$\epsilon^{2}$ ::: $i\partial_{t}(\bar{f}_{2}-f_{2})-\partial_{x}\partial_{t}(\bar{f}_{2}+f_{2})=\left[iD_{t}-D_{x}D_{t}\right](\bar{g}_{1}.g_{1}),$ (6.25)
$\epsilon^{4}$ ::: $i\partial_{t}(\bar{f}_{4}-f_{4})-\partial_{x}\partial_{t}(\bar{f}_{4}+f_{4})=\left[iD_{t}-D_{x}D_{t}\right](\bar{g}_{1}.g_{3}+\bar{g}_{3}.g_{1}-\bar{f}_{2}.f_{2}),$ (6.26)
@@
$\epsilon^{2n}$ ::: $i\partial_{t}(\bar{f}_{2n}-f_{2n})-\partial_{x}\partial_{t}(\bar{f}_{2n}+f_{2n})=$ (6.29)
@@
$\epsilon^{2}$ ::: $\partial_{x}(\bar{f}_{2}-f_{2})=-D_{x}(\bar{g}_{1}.g_{1}),$ (6.30)
$\epsilon^{4}$ ::: $\partial_{x}(\bar{f}_{4}-f_{4})=-D_{x}(\bar{g}_{1}.g_{3}+\bar{g}_{3}.g_{1}+\bar{f}_{2}.f_{2}),$ (6.31)
[…]
$\epsilon^{2n}$ ::: $\partial_{x}(\bar{f}_{2n}-f_{2n})=-D_{x}\left[\sum_{n_{1}+n_{2}=n-1}(\bar{g}_{2n_{1}+1}.g_{2n_{(}2+1)}+\sum_{n_{1}+n_{2}=n}\bar{f}_{2n_{1}}.f_{2n_{2}})\right].$ (6.35)
@@ 7 Nonlocal KE
The gauge equivalent spin system corresponding to the K-IIE is given by (1.3)-(1.4). But here we must note that in contrast to the local case, in our nonlocal case, in the Serret-Frenet equations (4.1), the curvatures $\kappa(t,x)$ and ${\bf\sigma}(t,x)$, the torsion $\tau(t,x)$, $\omega_{j}(t,x)$ are complex-valued functions. As results, in the nonlocal case, the spin matrix $S$ is not Hermitian and has $PT$ symmetry $S(t,x)=\sigma_{3}S^{+}(t,-x)\sigma_{3}$. The corresponding spin vector ${\bf S}(t,x)=(S_{1}(t,x),S_{2}(t,x),S_{3}(t,x))$ is complex-valued vector. As we mentioned above, in the nonlocal case, the spin matrix $S(t,x)$ is not Hermitian. But we can decompose it as the sum of a Hermitian matrix and a skew-Hermitain matrix as [39]
@@ 9.3 Multicomponent KE
where $k=1,2,...,N$. The 2-component version of this equation reads as

### v4

@@ 6.2 Hirota bilinear form of the K-IE
$\epsilon^{1}$ $:$ $ig_{1t}+g_{1xt}=0,$ (6.19)
$\epsilon^{3}$ $:$ $\left[i\partial_{t}+\partial_{x}\partial_{t}\right]g_{3}=\left[iD_{t}-D_{x}D_{t}\right](\bar{f}_{2}.g_{1}),$ (6.20)
$\cdot\cdot\cdot$ (6.21)
$\cdot\cdot\cdot$ (6.22)
$\cdot\cdot\cdot$ (6.23)
$\epsilon^{2n+1}$ $:$ $\left[i\partial_{t}+\partial_{x}\partial_{t}\right]g_{2n+1}=\sum_{k+m=n}\left[iD_{t}-D_{x}D_{t}\right](\bar{f}_{2k}.g_{2m+1}),$ (6.24)
@@
$\epsilon^{2}$ $:$ $i\partial_{t}(\bar{f}_{2}-f_{2})-\partial_{x}\partial_{t}(\bar{f}_{2}+f_{2})=\left[iD_{t}-D_{x}D_{t}\right](\bar{g}_{1}.g_{1}),$ (6.25)
$\epsilon^{4}$ $:$ $i\partial_{t}(\bar{f}_{4}-f_{4})-\partial_{x}\partial_{t}(\bar{f}_{4}+f_{4})=\left[iD_{t}-D_{x}D_{t}\right](\bar{g}_{1}.g_{3}+\bar{g}_{3}.g_{1}-\bar{f}_{2}.f_{2}),$ (6.26)
@@
$\epsilon^{2n}$ $:$ $i\partial_{t}(\bar{f}_{2n}-f_{2n})-\partial_{x}\partial_{t}(\bar{f}_{2n}+f_{2n})=$ (6.29)
@@
$\epsilon^{2}$ $:$ $\partial_{x}(\bar{f}_{2}-f_{2})=-D_{x}(\bar{g}_{1}.g_{1}),$ (6.30)
$\epsilon^{4}$ $:$ $\partial_{x}(\bar{f}_{4}-f_{4})=-D_{x}(\bar{g}_{1}.g_{3}+\bar{g}_{3}.g_{1}+\bar{f}_{2}.f_{2}),$ (6.31)
$\cdot\cdot\cdot$ (6.32)
$\cdot\cdot\cdot$ (6.33)
$\cdot\cdot\cdot$ (6.34)
$\epsilon^{2n}$ $:$ $\partial_{x}(\bar{f}_{2n}-f_{2n})=-D_{x}\left[\sum_{n_{1}+n_{2}=n-1}(\bar{g}_{2n_{1}+1}.g_{2n_{(}2+1)}+\sum_{n_{1}+n_{2}=n}\bar{f}_{2n_{1}}.f_{2n_{2}})\right].$ (6.35)
@@ 7 Nonlocal KE
The gauge equivalent spin system corresponding to the K-IIE is given by (1.3)-(1.4). But here we must note that in contrast to the local case, in our nonlocal case, in the Serret-Frenet equations (4.1), the curvatures $\kappa(t,x)$ and ${\bf\sigma}(t,x)$, the torsion $\tau(t,x)$, $\omega_{j}(t,x)$ are complex-valued functions. As results, in the nonlocal case, the spin matrix $S$ is not Hermitian and has $PT$ - symmetry $S(t,x)=\sigma_{3}S^{+}(t,-x)\sigma_{3}$. The corresponding spin vector ${\bf S}(t,x)=(S_{1}(t,x),S_{2}(t,x),S_{3}(t,x))$ is complex-valued vector. As we mentioned above, in the nonlocal case, the spin matrix $S(t,x)$ is not Hermitian. But we can decompose it as the sum of a Hermitian matrix and a skew-Hermitain matrix as [39]
@@ 9.3 Multicomponent KE
where k=1,2,,Nk=1,2,...,N. The 2-component version of this equation reads as
@@ 9.6 Integrable Nurshuak equation
10 The Akbota-Tolkynay-Zhaidary-Myrzakulov equation
Our next example of integrable equations is the Akbota-Tolkynay-Zhaidary-Myrzakulov equation (ATZME). The Akbota-Tolkynay-Zhaidary-Myrzakulov equation (ATZME) reads as:
$q_{t}-\frac{1}{b}uq_{x}+\frac{2\beta}{b}qq_{y}-\beta r_{y}$ $=$ $0,$ (10.1)
$r_{t}-\frac{1}{b}ur_{x}+\frac{2\beta}{b}rq_{y}-\frac{\beta}{2ab}q_{xxy}$ $=$ $0,$ (10.2)
$u_{x}+\beta q_{y}$ $=$ $0,$ (10.3)
where $a,b,\beta$ are real constants, $(q,r,u)$ are some functions of $(x,t,y)$. We note that the ATZME (10.1)-(10.3) is integrable. Its Lax equations looks like
$\Phi_{x}$ $=$ $U_{10}\Phi,$ (10.4)
$\Phi_{t}$ $=$ $\beta\lambda\Phi_{y}+B\Phi,$ (10.5)
where
$U_{10}$ $=$ $\begin{pmatrix}0&a\\ b\lambda^{2}+q\lambda+r&0\end{pmatrix}$, (10.6)
$B$ $=$ $B_{2}\lambda^{2}+B_{1}\lambda+B_{0},$ (10.7)
$B_{2}$ $=$ $\begin{pmatrix}0&0\\ u&0\end{pmatrix},\quad B_{1}=\begin{pmatrix}0&0\\ b^{-1}uq&0\end{pmatrix}$, (10.8)
$B_{0}$ $=$ $\begin{pmatrix}\frac{\beta}{2b}q_{y}&ab^{-1}u\\ b^{-1}ur+\frac{\beta}{2ab}q_{xy}&-\frac{\beta}{2b}q_{y}\end{pmatrix}$. (10.9)
The compatibility condition $\Phi_{xt}=\Phi_{tx}$ of the linear equations (10.4)-(10.5) that is
$U_{10t}-B_{x}+[U_{10},B]-\beta\lambda U_{10y}=0$ (10.10)
gives the ATZME (10.1)-(10.3). As the integrable equation, the ATZME (10.1)-(10.3) has the N-soliton solution, infinite number of conservation laws, Hamiltonian structure and so on.